I swing a ball around my head at constant speed in a circle with circumference 3 m. What is the work done on the ball by the 10 N tension force in the string during one revolution of the ball?

A.30 J

B.20 J

C.10 J

D.0 J

Why? I need help understanding this question. Should I use Krot=1/2Iw^2 equation???
Answer #1

W = Fs.

see work done is force dot product with distance

W=F.s = F*s*cos(theta)

the angle between tension force and distance is 90 degree

which implies

W=10*3*cos(90) = 0 joules